Eagle Company makes the MusicFinder, a sophisticated satellite radio. Eagle has experienced a steady growth in sales for the past five years. However, Ms. Luray, Eagle's CEO, believes that to maintain the company's present growth will require an aggressive advertising campaign next year. To prepare for the campaign, the company's accountant, Mr. Bednarik, has prepared and presented to Ms. Luray the following data for the current year, Year 1:


  Variable costs:
      Direct labor (per unit) $ 92  
      Direct materials (per unit) 43  
      Variable overhead (per unit) 14  
          Total variable costs (per unit) $ 149  
  
  Fixed costs (annual):
      Manufacturing $ 385,000  
      Selling 298,000  
      Administrative 799,000  
          Total fixed costs (annual) $ 1,482,000  
  Selling price (per unit) 412  
  Expected sales revenues, Year 1 (29,000 units) $ 11,948,000  


Eagle has an income tax rate of 40 percent.

     Ms. Luray has set the sales target for Year 2 at a level of $13,184,000 (or 32,000 radios).


Required:
(a)

What is the projected after-tax operating profit for Year 1?

(b)

What is the break-even point in units for Year 1? (Round up to the nearest whole unit.)

(c)

Ms. Luray believes that to attain the sales target (32,000 radios) will require additional selling expenses of $293,000 for advertising in Year 2, with all other costs remaining constant. What will be the after-tax operating profit for Year 2 if the firm spends the additional $293,000?

(d)

What will be the break-even point in sales dollars for Year 2 if the firm spends the additional $293,000 for advertising? (Solve by computing volume in units first. Round up to the nearest whole unit.)

(e)

If the firm spends the additional $293,000 for advertising in Year 2, what is the sales level in dollars required to equal the Year 1 after-tax operating profit? (Solve by computing volume in units first. Round up to the nearest whole unit.)

(f) At a sales level of 32,000 units, what is the maximum amount the firm can spend on advertising to earn an after-tax operating profit of $759,000?
Answer #1

a)

S. No. Particulars Amount
I Sales 11948000
II less; Total variable cost(149*29000) (4321000)
III Contribution (I-II) 7627000
IV less; Total Fixed cost (1482000)
V profit before tax(III-IV) 6145000
VI less; Income tax @40% (2458000)
VII Profit after tax $3687000

b) VCR ratio = total variable cost per unit / selling price per unit*100

= 149/412*100 =36.17 %

hence, PVR = 63.83% (1- 36.17%)

at the break-even point, PVR = Fixed cost/ sales(volume)

so 63.83% = 1482000/ break even sales

break even sales = 1482000/63.83% = $2321792

c)

S. No. Particulars Amount
I Sales (412*32000) 13184000
II less; Total variable cost(149*32000) 4768000
III Contribution (I-II) 8416000
IV less; Total Fixed cost(1482000+293000) 1775000
V profit before tax(III-IV) 6641000
VI less; Income tax @40% 2656400
VII Profit after tax 3984600

d) break-even point sales (in units) = Fixed cost/contribution per unit

= 1775000/(412-149)=1775000/263 = 6749

e)sales level in dollars required to equal the Year 1 after-tax operating profit

= profit before tax of year I +fixed cost of year II should be equal to the contribution

sales at units = 6145000+1775000/contribtuion per unit

= 7920000/263 =30114

f) at sales 32000

S.No. Particulars Amount
I sales (32000*412) 13184000
II contribution(32000*263) 8416000
III less; profit required (759,000)
IV Total Fixed cost(II-III) 7657000
V Fixed cost excluding advertisement (1482000)
VI AllowableAdvertisement expenses (IV-V) 6175000